OpenStudy (anonymous):

can someone help me with a few questions. v=p(1+rt) p=$500 r=.08 per year t=3 years s=4.9t2 find s if t- 4 seconds p=l2r l= 12 amperes r= 2ohms a=2w2+4hw find a if w=7cm h=10cm

OpenStudy (anonymous):

for the top one its $620

OpenStudy (anonymous):

the second one: s = 4.9t^2 s = 4.9(4^2) s = 4.9(16) s = 78.4

OpenStudy (anonymous):

I don't understand what the other questions are asking.

OpenStudy (anonymous):

p=l^2r find p if l=12 amperes r-2 ohms a=2w^2 + 4hw find a if w=7cm h=10cm